When I was 4, my brother was half my age. Now I am 100, how old is he
Found 2 solutions by prashanthi, Maths68:
Answer by prashanthi(2)

when i was 4
my brother was 1/2 my age i.e 2
when i am 100
he is 1/2(100)=50

Answer by Maths68(1474)

You were = 4 years
You brother was = 2 years (half of your age)
You are = (4+96) =100 years old now
Your brother is = (2+96) = 98 years old now